# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v1

@@ Abstract.
Efficient use of GPU memory is essential for high throughput LLM inference. Prior systems reserved memory for the KV-cache ahead-of-time, resulting in wasted capacity due to internal fragmentation. Inspired by OS-based virtual memory systems, vLLM proposed PagedAttention to enable dynamic memory allocation for KV-cache. This approach eliminates fragmentation, enabling high-throughput LLM serving with larger batch sizes. However, to be able to allocate physical memory dynamically, PagedAttention changes the layout of KV-cache from contiguous virtual memory to non-contiguous virtual memory. This change requires attention kernels to be rewritten to support paging, and serving framework to implement a memory manager. Thus, the PagedAttention model leads to software complexity, portability issues, redundancy and inefficiency.
In this paper, we propose vAttention for dynamic KV-cache memory management. In contrast to PagedAttention, vAttention retains KV-cache in contiguous virtual memory and leverages low-level system support for demand paging, that already exists, to enable on-demand physical memory allocation. Thus, vAttention unburdens the attention kernel developer from having to explicitly support paging and avoids re-implementation of memory management in the serving framework. We show that vAttention enables seamless dynamic memory management for unchanged implementations of various attention kernels. vAttention also generates tokens up to 1.97×1.97$\times$1.97 × faster than vLLM, while processing input prompts up to 3.92×3.92$\times$3.92 × and 1.45×1.45$\times$1.45 × faster than the PagedAttention variants of FlashAttention and FlashInfer.
copyright: none
@@ 1. Introduction
Large Language Models (LLMs) are being deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (OpenAI, 2023; Chowdhery et al., 2022; cla, [n. d.]; bin, [n. d.]; bar, [n. d.]; git, [n. d.]; ama, [n. d.]; rep, [n. d.]). Optimizing LLM inference has thus become important (Agrawal et al., 2023; Holmes et al., 2024; Patel et al., 2023; Zhong et al., 2024; Hu et al., 2024; Kwon et al., 2023; Yu et al., 2022). One of the key techniques used to improve LLM serving throughput is batching (Yu et al., 2022; Kwon et al., 2023; Patel et al., 2023; Agrawal et al., 2024). Out of the two phases of LLM inference – namely prefill and decode – the decode phase is memory-bound because it processes a single token at-a-time per request. Batching amortizes the cost of fetching model weights from GPU memory and boosts throughput by improving memory bandwidth utilization (Agrawal et al., 2023).
@@
vLLM: Pioneered PagedAttention. Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.85×2.85$\times$2.85 × slower than FlashAttention (9(b)). Further, block size has as much as 3×3$\times$3 × effect on the execution time of the PagedAttention kernel (Figure 5).
@@
FlashAttention: PagedAttention version of the kernel is up to 12% slower than the vanilla version. Initial attempts to add paging support failed unit tests (fa-, 2023).
@@
TensorRT-LLM: Serving throughput dropped by more than 10% in a Python front-end (ten, 2023). Recommends using the C++ frontend. Even with C++, we observe up to 5% higher latency in some cases with PagedAttention.
FlashInfer: Using PagedAttention-based prefill kernel increases time-to-first-token (TTFT) by up to 45% (Table 6).
@@
Table 1. The PagedAttention model requires an application to explicitly manage dynamically allocated physical memory. These examples highlight the complexity, performance and maintenance challenges associated with this approach.
Efficient inference also requires careful allocation of GPU memory. For every request, an LLM maintains an in-memory state known as the KV-cache and re-uses it in every iteration for the lifetime of the request (Yu et al., 2022; Agrawal et al., 2023; Patel et al., 2023). Achieving high memory capacity utilization during inference is challenging for two reasons: 1) per-request KV-cache grows slowly i.e., one token per iteration (few 10s of milliseconds) and 2) the number of generated tokens, and hence the total size of a request’s KV-cache, is typically not known ahead-of-time.
Prior systems like Orca (Yu et al., 2022) and FasterTransformer (fas, 2024b) allocate a contiguous chunk of virtual memory (backed by pre-allocated physical memory) for the KV-cache. The allocated size corresponded to model’s maximum supported sequence length, e.g., 32K. Since models often generate far fewer tokens than the maximum limit, significant GPU memory was wasted due to internal fragmentation. Consequently, these systems exhibit poor throughput as they are unable to support large batch sizes.
System Memory management in user code KV-cache fragmentation KV-cache allocation Portable wrt GPU code
Programming Effort
GPU CPU
vLLM (Kwon et al., 2023) Block-Table(s) Low Fast No High High
FlashInfer (Ye et al., 2024) Compressed Block-Table(s) Low Fast No High High
vAttention None Low Slow Yes None Low
Table 2. Dynamic memory management in various systems. TensorRT-LLM (trt, 2023), LightLLM (lig, 2023) and FlashAttention (Dao, 2023) also adopt vLLM-style PagedAttention model. FlashInfer uses a compressed representation and prefetching to limit the overhead of paging. We introduce various LLM-specific optimizations to hide the impact of high allocation latency in vAttention.
@@
Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (Kwon et al., 2023) to mitigate KV-cache related memory fragmentation. Instead of reserving the maximum sequence length of KV-cache memory ahead-of-time, vLLM allocates small blocks of virtual memory (backed by physical memory) on-demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens. However, dynamically allocated blocks are not guaranteed to be contiguous in virtual memory (the system may have allocated those to other requests). Thus, PagedAttention accepts that KV-cache allocations will be non-contiguous and implements a block-table to stitch together these non-contiguous allocations (§3.2).
Today, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems e.g., in TensorRT-LLM (trt, 2023), HuggingFace TGI (hft, [n. d.]), FlashInfer (Ye et al., 2024), LightLLM (lig, 2023) etc. The most notable aspect of the PagedAttention approach is that it stores KV-cache in non-contiguous virtual memory to be able to allocate physical memory dynamically. While this approach provides an adequate solution for KV-cache fragmentation, we argue that it has several pitfalls (see Table 1 for empirical evidence and real-world experiences):
@@
1. Requires re-writing the attention kernel (GPU code). The elements of a virtually contiguous object can be accessed using index-based lookup which is both simple and efficient [^1] A popular example of this is arrays vs. linked lists.. By storing KV-cache in non-contiguous virtual memory, PagedAttention mandates re-writing GPU code so that the attention kernel can de-reference all the elements of KV-cache. The need to re-write code is a major barrier to using new attention optimizations in production settings.
@@
2. Adds software complexity and redundancy (CPU code). PagedAttention also forces the developers to implement a memory manager inside the serving framework, making it responsible for (de)allocating KV-cache and tracking the location of dynamically allocated KV-cache blocks. This approach essentially translates to re-implementing demand paging – which is an OS functionality – in user code.
@@
3. Introduces performance overhead. PagedAttention can add runtime overhead in the critical path of execution in two ways. First, it requires GPU kernels to execute extra code related to fetching KV-cache from non-contiguous memory blocks. We show that this can slow down attention computation by more than 10% in many cases. Second, the user space memory manager can add CPU overhead, contributing up to another 10% cost (§3.3).
Figure 1. Illustration of memory waste due to internal fragmentation. Orca (top) reserves memory for the KV-cache ahead-of-time. vLLM (bottom) mitigates fragmentation with dynamic memory allocation. Shaded boxes represent memory occupied by the KV tokens whereas the white boxes represent allocated but unused memory.
In this paper, we argue that retaining the virtual memory contiguity of KV-cache is critical for reducing software complexity and redundancy in LLM deployments. Instead of re-implementing paging at the user-level, we contend that existing virtual memory abstractions in the OS can be re-purposed for dynamic KV-cache memory management, resulting in simplified deployments as well as higher performance.
@@
To demonstrate this, we propose vAttention – a system that stores KV-cache in contiguous virtual memory without committing physical memory ahead-of-time. We achieve this by leveraging CUDA support of low-level virtual memory APIs which expose distinct interfaces for allocating virtual and physical memory (§5). vAttention exposes a set of simple APIs using which a serving framework reserves contiguous virtual space for the KV-cache and allocates physical memory on-demand. This approach lends several benefits as listed in Table 2. vAttention also improves portability by enabling a seamless re-use of readily available GPU kernels while eliminating the need to implement a memory manager in a serving system.
@@
Challenges and Optimizations: vAttention solves two key challenges in enabling efficient dynamic memory management without PagedAttention (§6). First, the minimum physical memory allocation granularity supported by CUDA APIs is 2MB. This size can result in significant wasted capacity depending on the model and workload characteristics (Table 8). To address this, we modify the open-source CUDA unified virtual memory driver to add support for finer-grained physical memory allocations of 64KB to 256KB. Second, memory allocation using CUDA APIs incurs high latency because each allocation involves a round-trip to the kernel. To hide the latency of memory allocation from end-users, we introduce several LLM-specific optimizations such as overlapping memory allocation with compute, executing some operations ahead of time and deferring memory reclamation. We show that our optimizations make vAttention an efficient KV-cache memory manager.
@@
  • We present vAttention, a system that retains the virtual contiguity of KV-cache while enabling dynamic allocation of physical memory.
  • We show that vAttention is able to seamlessly add dynamic memory management support to unmodified attention kernels of FlashAttention (fag, 2022) and FlashInfer (fig, 2023) while also being performant.
  • We evaluate Yi-6B, Llama-3-8B and Yi-34B on 1-2 A100 GPUs and show that using FlashAttention’s original kernel, vAttention outperforms vLLM by up to 1.97×1.97$\times$1.97 ×, while reducing the time-to-first-token (TTFT) by up to 1.45×1.45$\times$1.45 × over the PagedAttention variant of FlashInfer.
@@ 2.1. Large Language Models
Given an input sequence, an LLM predicts the probability of an output sequence wherein a sequence is a set of tokens (Kwon et al., 2023). Each inference request begins with a prefill phase that processes all its prompt tokens in parallel. The prefill phase produces the first output token of a request. Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (Agrawal et al., 2023).
LLMs are built atop one of variants of the transformer architecture (Vaswani et al., 2017). A transformer block contains two types of operators: position-wise and sequence-wise. The former category includes feed-forward network, layer normalization, activation, embedding layer, output sampling layer, and residual connections whereas attention is a sequence-level operator. In this paper, we primarily focus on attention since it is the primary consumer of GPU memory in LLM inference.
@@
The attention score is computed separately for each request in the batch. Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention [^2] An exception to this is sliding window attention that fixes an upper bound on how many of the prior tokens are used in computing attention.. Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations: we refer to this state as KV-cache. A request is executed until the model generates a special end-of-sequence token or reaches the maximum context length for the request.
Structure of the KV-cache and terminology: An LLM consists of multiple layers of the transformer block and each layer maintains its own cache of keys and values. In this paper, we refer to the cache of all transformer blocks collectively as KV-cache while using the term K-cache or V-cache for keys and values, respectively. In deep learning frameworks, the K-cache (or V-cache) at each layer is typically represented as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request. We refer to a kernel implementation that computes attention scores over contiguously stored K and V as a vanilla kernel.
@@ 2.2. Fragmentation and PagedAttention
To improve serving throughput, production systems rely on batching which requires careful allocation of GPU memory. This is challenging because the total context length of a request is not known in advance. Serving systems worked around this challenge by pre-reserving KV-cache space assuming that each context is as long as the maximum length supported by the model (e.g.,200K for Yi-6B-200K). vLLM shows that this strategy is prone to severe internal fragmentation. In fact, vLLM showed that prior reservation wastes memory even if the context lengths are known in advance. This is because the per-request KV-cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
Inspired by the OS-based virtual memory systems, vLLM proposed PagedAttention to mitigate fragmentation by dynamically allocating memory for the KV-cache. PagedAttention splits KV-cache into fixed-sized blocks and allocates memory for one block at a time. This way, vLLM allocates only as much memory as a request needs, and only when required – not ahead-of-time. Figure 1 shows an example of how reservation-based systems such as Orca (Yu et al., 2022) can waste significant memory due to fragmentation and how vLLM avoids it with dynamic memory allocation.
Figure 2. PagedAttention involves two layers of address translation. First, the framework tracks the virtual memory addresses of KV-cache blocks via Block-Tables. Second, the GPU performs virtual-to-physical address translation via OS-managed Page-Tables.
@@ 3. Issues with the PagedAttention Model
Despite being inspired by demand paging, PagedAttention adopts an approach that is different from conventional demand paging: it requires an application’s code to be modified to adapt to dynamically allocated physical memory whereas conventional demand paging is transparent to applications. This section elaborates on some of the issues that arise with such an approach.
@@ 3.1. Requires re-writing the attention kernel
PagedAttention necessitates re-writing the attention kernel. This is because conventional implementations of the attention operator assume that the two input tensors K and V (Equation 2) are stored in contiguous memory. By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV-cache blocks. Writing correct and performant GPU kernels can be challenging for most programmers (fa-, 2023).
Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (Dao et al., 2022; Dao, 2023; fla, 2023; Hong et al., 2024; Ye et al., 2024; Zhang et al., 2023; Bikshandi and Shah, 2023; Ainslie et al., 2023; Shazeer, 2019; Child et al., 2019; Beltagy et al., 2020), and this trend is likely to continue. In the PagedAttention model, keeping up with new research requires continued efforts in porting new optimizations to a PagedAttention-aware implementation. Production systems can therefore easily fall behind research, potentially losing performance and competitive advantage. To provide an example, 9(b) shows that the paged kernel of vLLM is already up to 2.85×2.85$\times$2.85 × slower than the FlashAttention counterpart for grouped-query attention (Ainslie et al., 2023).
Figure 3. Overhead of PagedAttention in FlashAttention’s decode kernel (model: Yi-6B). Legend shows the batch size.

### v2

@@ Abstract.
Efficient management of GPU memory is essential for high throughput LLM inference. Prior systems used to reserve KV-cache memory ahead-of-time that resulted in wasted capacity due to internal fragmentation. Inspired by demand paging, vLLM proposed PagedAttention to enable dynamic memory allocation for KV-cache. This approach eliminates fragmentation and improves serving throughout. However, to be able to allocate physical memory dynamically, PagedAttention changes the layout of KV-cache from contiguous virtual memory to non-contiguous virtual memory. As a consequence, one needs to rewrite the attention kernels to support paging, and implement a memory manager in the serving framework. This results in both performance and programming overheads, as well as portability challenges in adopting state-of-the-art attention kernels.
In this paper, we propose vAttention, a new approach for dynamic KV-cache memory management. In contrast to PagedAttention, vAttention stores KV-cache in contiguous virtual memory and leverages OS support for on-demand allocation of physical memory. vAttention thus enables one to use state-of-the art attention kernels out-of-the-box by adding support for dynamic allocation of physical memory without having to re-write their code. We implement vAttention in the vLLM serving stack to show that it also helps improve decode throughput by up to 1.99×1.$9$ over vLLM, and the end-to-end serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
@@ 1. Introduction
Large Language Models (LLMs) are deployed in a wide range of applications e.g., chat bots, search engines and coding assistants (openai2022gpt4techreport, ; chowdhery2022Palm, ; bingai, ; bard, ; githubcopilot, ; amazoncodewhisperer, ; replitghostwriter, ). Given the size and scale of modern LLM deployments, optimizing inference has become extremely important  (sarathi2023, ; splitfuse2024, ; patel2023splitwise, ; distserve2024, ; tetriinfer, ; vllmsosp, ; orca, ). LLM inference happens in two phases: a compute-bound prefill phase, followed by several iterations of memory-bound decode phase. Batching is a powerful technique to boost LLM serving throughput (orca, ; vllmsosp, ; patel2023splitwise, ; sarathiserve2024, ), as it amortizes the cost of fetching model weights from HBM to the on-device SRAM during decodes.
@@
vLLM: Pioneered PagedAttention. Despite being in an actively maintained code repository, vLLM’s PagedAttention kernel is up to 2.8×2.8$\times$2.8 × slower than FlashAttention (Table 6). Further, changing block size changes the execute latency of the kernel by as much as 3×3$\times$3 × (Figure 3).
@@
FlashAttention: PagedAttention based prefill kernel is up to $28\%$ slower than the non-paged kernel (Figure 2) while the decode kernel is up to $12\%$ slower. Initial attempts to add paging support failed unit tests (fa-paged-crash, ).
@@
TensorRT-LLM: Serving throughput dropped by more than $10\%$ in a Python front-end (tensorrt-perf-decay, ). Recommends using the C++ front-end. Even with C++, we observe up to $5\%$ higher latency in some cases with PagedAttention.
FlashInfer: PagedAttention based prefill kernel is up to $24\%$ slower than the non-paged kernel (Figure 2).
@@
Table 1. The PagedAttention approach requires an application to explicitly manage dynamically allocated physical memory, including re-writing of attention kernels. These examples highlight the complexity, performance and maintenance challenges associated with this approach.
Careful allocation of GPU memory is key to enabling larger batch sizes. For each request, the serving framework stores the activations of all the tokens processed so far in GPU memory and reuses them for the subsequent token generation. This is called the KV-cache (orca, ; sarathi2023, ; patel2023splitwise, ) which accounts for a majority of GPU memory usage during inference. Efficiently allocating GPU memory for the KV-cache of a batch of requests is challenging for two reasons. First, the per-request KV-cache grows slowly (one token per iteration), and second, a request’s decode length (or its total KV-cache size) is not known ahead of time.
Prior systems like Orca (orca, ) and FasterTransformer (fastertransformer, ) allocate GPU memory for each request based on the maximum context length supported by the model (e.g., Yi-34B model supports context length of up to 200K (yi-34b-200k-hf, )). However, in practice, the number of decode tokens generated are far less; the average decode length for the chat-based sharegpt dataset is 415 tokens (sarathiserve2024, ). Therefore, static memory allocation results in severe internal fragmentation, limiting batch size and serving throughput.
@@
Inspired by demand paging in OS-based virtual memory systems, vLLM introduced PagedAttention (vllmsosp, ) to mitigate KV-cache related memory fragmentation. Instead of reserving the maximum sequence length of KV-cache memory ahead-of-time, vLLM allocates small blocks of GPU memory on demand i.e., when previously allocated blocks are fully utilized and the model continues to generate more tokens. This approach provides a near-perfect solution for mitigating fragmentation and hence, PagedAttention has become the de facto standard for dynamic memory allocation in LLM serving systems e.g., in TensorRT-LLM (trtllmgithub, ), HuggingFace TGI (hftgi, ), LightLLM (lightllm:github, ) etc.
However, the PagedAttention approach faces a fundamental consequence of dynamic memory allocation: dynamically allocated memory blocks are not guaranteed to be contiguous. Note that user-level objects are allocated in virtual memory. Therefore, in trying to enable dynamic allocation of physical memory, PagedAttention ends up changing the layout of KV-cache from contiguous to non-contiguous virtual memory. We argue that this approach has several pitfalls (see Table 1 for empirical evidence and real-world experiences):
@@
1. Requires re-writing the attention kernel (GPU code). The elements of a virtually contiguous object can be accessed using index-based lookup which is both simple and efficient. Many implementations of deep learning operators also expect virtual contiguity of input tensors. However, by storing KV-cache in non-contiguous virtual memory, PagedAttention mandates re-writing GPU code so that the attention kernel can de-reference all the elements of KV-cache. The need to re-write code is a major barrier to using new attention optimizations in production settings.
@@
2. Adds software complexity and redundancy (CPU code). To stitch together dynamically allocated virtual memory blocks, PagedAttention forces developers to implement a memory manager in the serving framework. Typically, this requires a Block-Table manager in user space which is essentially a re-implementation of demand paging (an OS functionality) in user code.
@@
3. Introduces performance overhead. PagedAttention adds runtime overhead in the critical path of execution in two ways. First, GPU kernels that compute attention over a non-contiguous KV-cache need to execute additional instructions. In many cases, this slows down an attention kernel by more than 20%, compared to when the KV-cache is virtually contiguous. Second, the user space memory manager can add CPU overhead contributing up to another 10% cost (§3.3).
In this paper, we instead advocate for leveraging the OS support of virtual memory and demand paging for dynamic KV-cache memory management. This approach alleviates the aforementioned issues that appear with PagedAttention.
@@
To support our claims, we present the design and implementation of vAttention. Leveraging OS support, vAttention stores KV-cache in contiguous virtual memory without committing physical memory ahead-of-time, thereby enabling the use of state-of-the art high-performant attention kernels out-of-the box. We achieve this by using CUDA support of low-level virtual memory APIs which expose distinct interfaces for allocating virtual and physical memory (§5).
@@
Leveraging OS support for memory allocation in an LLM serving system poses two key efficiency challenges (§5.4). First, the minimum physical memory allocation granularity supported by CUDA is 2MB. This can result in significant wasted capacity and fragmentation. We address this challenge by modifying the open-source CUDA unified virtual memory driver, adding support for finer-grained physical memory allocations in multiples of 64KB. Our evaluation shows that using 64KB pages as the granularity of virtual-to-physical address translation does not affect the execution latency of attention kernels i.e., we do not find any evidence of TLB thrashing. Second, memory allocation using CUDA APIs incurs high latency because each allocation involves a round-trip to the OS kernel. We introduce several LLM-specific optimizations such as overlapping memory allocation with compute, opportunistically allocating pages ahead of time, and deferring memory reclamation. These optimizations hide the latency cost of memory allocation from end-users, making vAttention an efficient KV-cache memory manager.
@@
  • We present vAttention – a memory management approach that retains the virtual contiguity of KV-cache while enabling dynamic physical memory allocation.
  • We implement vAttention in vLLM serving stack to show that it seamlessly adds dynamic memory management support to unmodified attention kernels of FlashAttention (fagithub, ) and FlashInfer (figithub, ).
  • We compare vAttention against PagedAttention based alternatives of vLLM, FlashAttention and FlashInfer by evaluating Yi-6B, Llama-3-8B and Yi-34B on 1-2 A100 GPUs. Using FlashAttention’s non-paged kernel, vAttention outperforms vLLM by up to 1.99×1.$9$ in decode throughput, while improving the end-to-end LLM serving throughput by up to 1.22×1.$2$ and 1.29×1.29$\times$1.29 ×, compared to using the state-of-the-art PagedAttention based kernels of FlashAttention and FlashInfer.
@@ 2.1. Large Language Models
Given an input sequence, an LLM predicts the probability of an output sequence wherein a sequence is a set of tokens (vllmsosp, ). Each inference request begins with a prefill phase that processes all its prompt tokens in parallel. The prefill phase produces the first output token of a request. Thereafter, the decode phase iteratively processes the output token generated in the previous step and produces the next output token in every iteration (sarathi2023, ).
LLMs are built atop one of variants of the transformer architecture (attentionpaper, ). A transformer block contains two types of operators: position-wise and sequence-wise. The former category includes feed-forward network, layer normalization, activation, embedding layer, output sampling layer, and residual connections whereas attention is a sequence-level operator. In this paper, we primarily focus on attention since it is the primary consumer of GPU memory in LLM inference.
@@
The attention score is computed separately for each request in the batch. Note that in each iteration of a request, all its preceding $k_{i}$ and $v_{i}$ are needed to compute attention. Hence, an inference engine stores the $k_{i}$ and $v_{i}$ vectors in memory to reuse them across iterations: we refer to this state as KV-cache. A request is executed until the model generates a special end-of-sequence token or reaches the maximum context length for the request.
Structure of the KV-cache and terminology: An LLM consists of multiple layers of the transformer block and each layer maintains its own cache of keys and values. In this paper, we refer to the cache of all transformer blocks collectively as KV-cache while using the term K-cache or V-cache for keys and values, respectively. In deep learning frameworks, the K-cache (or V-cache) at each layer is typically represented as a 4D tensor of shape $[B,L,H,D]$ where $B$ refers to batch size and $L$ refers to the maximum possible context length of a request. We refer to a kernel implementation that computes attention scores over contiguously stored K and V as a non-paged kernel.
@@ 2.2. Fragmentation and PagedAttention
To improve serving throughput, production systems rely on batching which requires careful allocation of GPU memory. This is challenging because the total context length of a request is not known in advance. Serving systems worked around this challenge by pre-reserving KV-cache space assuming that each context is as long as the maximum length supported by the model (e.g., 200K for Yi-34B-200K). vLLM shows that this strategy is prone to severe internal fragmentation. In fact, vLLM showed that prior reservation is suboptimal even if the context lengths are known in advance. This is because the per-request KV-cache grows one token at a time and hence prior reservation wastes memory over the entire lifetime of a request.
Inspired by the OS-based virtual memory systems, vLLM proposed PagedAttention to mitigate fragmentation by dynamically allocating memory for the KV-cache. PagedAttention splits KV-cache into fixed-sized blocks and allocates memory for one block at a time. This way, vLLM allocates only as much memory as a request needs, and only when required – not ahead-of-time.
Figure 1. PagedAttention involves two layers of memory management: one in user space and one in OS kernel space.
@@ 3. Issues with the PagedAttention Model
Despite being inspired by demand paging, the PagedAttention approach is different from it: PagedAttention mandates an implementation of paging in user space whereas conventional demand paging is transparent to applications. This section elaborates on issues that arise with such an approach.
@@ 3.1. Requires Re-writing the Attention Kernel
PagedAttention necessitates re-writing the attention kernel. This is because conventional implementations of the attention operator assume that the two input tensors K and V (Equation 2) are stored in contiguous memory. By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV-cache blocks. Writing correct and performant GPU kernels can be challenging for most programmers (fa-paged-crash, ).
Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (flashattention, ; flashattention2, ; flashdecoding, ; flashdecoding++, ; flashinfer, ; h2o2023, ; flashattnhopper, ; groupedqueryattention, ; multiqueryattention, ; generatinglongsequence2019, ; longformer, ), and this trend is likely to continue. In the PagedAttention model, keeping up with new research requires continued efforts in porting new optimizations to a PagedAttention-aware implementation. Production systems can therefore easily fall behind research, potentially losing performance and competitive advantage. To provide an example, Table 6 shows that the paged kernel of vLLM is already up to 2.8×2.8$\times$2.8 × slower than the FlashAttention kernel (groupedqueryattention, ).